ou are given the data below for 2008 for the imaginary country of Amagre, whose currency is the G.


Consumption 350 billion G

Transfer payments 100 billion G

Investment 100 billion G

Government purchases 200 billion G

Exports 50 billion G

Imports 150 billion G

Bond purchases 200 billion G

Earnings on foreign investments 75 billion G

Foreign earnings on Amagre investment 25 billion G

Compute net foreign investment.

Compute net exports.

Compute GDP.

Compute GNP.

In addition to responding with a quantitative answer, briefly describe how you arrived at your answers.


1
Expert's answer
2021-10-13T10:17:55-0400

Net foreign investment = Earning on foreign investment - foreign earnings on amagre investment 


a. Net foreign investment = 75 billion - 25 billion

  Net foreign investment = 50 billion

 

b) Net exports = Exports - Imports

   Net export = 50 - 150 billion

Net export = - 100 billion

 


C) GDP = C + I + G + (X-M)

   GDP = 300 + 100 + 200 + (50 - 150)

   GDP = 600 - 100

GDP = 500 billion G

 


d) GNP = GDP + Net foreign Investment

   GNP = 500 + 50 billion G

GNP = 550 billion G